OpenStudy (anonymous):

Express x/3-1/x^2/3-3 in simplest form.

OpenStudy (anonymous):

\[\frac{ x }{ 3 }-1\div \frac{ x ^{2} }{ 3 }-3\]

OpenStudy (tkhunny):

Still no. Use parentheses. Like all division of fractions, try the old "Reciprocal and Multiply" idea. Of course, you may have to do this first: \(\dfrac{x}{3} - 1 = \dfrac{x-3}{3}\)

OpenStudy (anonymous):

@tkhunny. So then x=3if I'm right?

OpenStudy (campbell_st):

well you have an expression so its no use solving it... the suggestion given is to get all terms over a common denominator so \[\frac{x}{3} - 1 = \frac{x}{3} - \frac{3}{3} = \frac{x - 3}{3}\] the difficulty is understanding your question is it \[(\frac{x}{3} -1) \div (\frac{x^2}{3} - 3)\] until that is decided, its difficult to givn you advice on how to simplify this problem

OpenStudy (anonymous):

@campbell_st yes that is the question

OpenStudy (campbell_st):

ok... now you need a common denominator for the 2nd term \[\frac{x^2}{3} - 3 = \frac{x^2}{3} - \frac{9}{3} = \frac{x^2 - 9}{3}\] so your problem becomes \[\frac{x -3}{3} \div \frac{x^2 - 9}{3}\] the rule for dividing by a fraction.. in simple terms is flip and multiply so the problem becomes \[\frac{x -3}{3} \times \frac{3}{x^2 - 9}\] hopefully you can work from here... and you may need to factor the new denominator to simplify the fraction.

OpenStudy (anonymous):

@campbell_st So after i do that it would look like this \[\frac{ x-3}{ 3 } * \frac{ 3 }{ x ^{2}-9 }\] and after I do that it would lead to this \[3*\frac{ 3 }{ x+3 }\rightarrow \frac{ 9 }{ x+3}\] Would this be right?

OpenStudy (campbell_st):

not quite the 3's will cancel and you are left with \[\frac{x - 3}{3} \times \frac{3}{x^2 - 9} = \frac{x - 3}{x^2 - 9}\] as I said before, factor the denominator and you will find another common factor that can be removed

OpenStudy (anonymous):

So then it would be \[\frac{ x+3 }{ x-3 }\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

not quite, what is the factored form of \[x^2 - 9\]

OpenStudy (anonymous):

\[(x-3) (x+3)\]

OpenStudy (campbell_st):

great so your problem is now \[\frac{(x -3)}{x^2-9} = \frac{(x -3)}{(x -3)(x+3)}\] now remove the common factor for your answer

OpenStudy (anonymous):

oh ok so then the answer would be (x+3)

OpenStudy (campbell_st):

no... remember it's a fraction, it needs a numerator given the denominator is (x + 3)